OpenStudy (anonymous):

In the below system, solve for y in the first equation. 2x + y = 11 5x – 2y = –1 a.2x + 11 b.–2x + 11 c.2x – 11 d.–2x – 11

OpenStudy (anonymous):

2x + y = 11 2x -2x + y = 11 -2x We want to cancel the 2x from the left side so we subtract 2x to both sides.

OpenStudy (anonymous):

y = -2x +11